# Mahyar Habibi,1 Dirk Hovy,2 Carlo Schwarz1

### v1

@@ 2.2 Measuring Distortions to Online Content
As described in the introduction, our measure of content distortions is built around the idea of the semantic space. To build intuition, imagine that the semantic content of a text can be represented as a vector in potentially infinite-dimensional semantic space. In this space, texts that talk about the same issue or hold the same opinion are close together, while texts about other issues or diverging opinions are far apart. Our measure of content distortions will measure to what extent the removal of toxic content changes the mean and the variance of the semantic space. The measure will be small if the semantic space after the removal of toxic content is very similar to the original semantic space prior to content moderation. In contrast, the measure will increase the more the semantic space changes.
Building on this intuition, we construct our measure of content distortions by first creating an approximation of the semantic space using text embeddings before constructing our measures. We describe each of these steps in the following. First, we create embeddings for each of the Tweets in our data using the BERTweet model (Nguyen et al., 2020). [^9] We also provide robustness checks for embeddings created by the RoBerta and DeBerta models. The BERTweet model, trained on a large English-language Twitter corpus, transforms the text of each Tweet into a 768-dimensional vector. In this way, the BERTweet model generates an approximation of the semantic space. [^10] Additional details on the BERTweet model can be found in Section B.1.
@@
As the second step, we construct a measure of content distortion based on the embeddings matrix $\mathbf{X}$. Our measure is based on the Bhattacharyya distance (BCD), a commonly used metric for comparing probability distributions. As the embeddings are approximately normally distributed (see Appendix Figure B.1), we use the formula for Bhattacharyya distance for multivariate normal distributions (e.g., Abou-Moustafa and Ferrie, 2012). For the case of two multivariate normal distributions $\mathcal{N}_{1}(\mu_{1},\Sigma_{1})$ and $\mathcal{N}_{2}(\mu_{2},\Sigma_{2})$, the BCD is defined as:
@@
where $\Sigma=\frac{\Sigma_{1}+\Sigma_{2}}{2}$. The Bhattacharyya intuitively consists of two additive terms. The first term is the squared Mahalanobis distance (Mahalanobis, 1936). It measures the inverse-variance-weighted distortions to the mean of the multivariate normal distribution, while the second term measures distortions to the variance. The $\det\Sigma$ is the so-called generalized variance index (GVI) (Wilks, 1932), which provides a multivariate extension of the standard statistical variance measure based on the mean squared deviation.
To calculate the BCD, we calculate both the means ($\mu$) and the variance-covariance matrix ($\Sigma$) of the embedding matrix $\mathbf{X}$ [^11] As the calculation of the dot product $XX^{\prime}$ is computationally unfeasible, we instead use a maximum likelihood estimator of the empirical covariance matrix using the procedure implemented in scikit learn. In tests on a subsample of the data, we confirmed that the approximate is very close to the analytical solution of the variance-covariance matrix. The diagonal elements of the variance-covariance matrix capture the dispersion of Tweets along the embedding dimensions. Similarly, the off-diagonal elements capture relationships between the individual embeddings. [^12] The individual entries in $\Sigma(\mathbf{X})$ are relatively small, which can lead to integer underflow when calculating the GVI. To avoid this issue, we multiply the embedding matrix $\mathbf{X}$ by 100 before calculating the BCD. Given that this equally affects all components of the BCD, this has no bearing on our results.
In this way, BCD provides a unidimensional measure that summarizes the overall distortions of the embeddings matrix $\mathbf{X}$. The BCD has several properties that make it a desirable measure for our application:
@@
    Another key advantage of the BCD is its computational tractability in high-dimensional settings. The BCD allows us to approximate the distortions of the embedding space in a computationally feasible manner, a crucial aspect given the expansive nature of online platforms and our data. [^13] Many other alternative methods, such as the ones that involve the computation of convex hulls, become computationally infeasible in high dimensions settings.
@@ 3 Results
In the following, we use the BCD to establish four sets of results. First, we analyze how the content is distorted when we remove highly toxic Tweets from the data. This analysis allows us to simulate the effect of more stringent content moderation. Second, we demonstrate that the content moderation-induced distortions are not driven by the toxic language of the Tweets. Third, we investigate what underlies the changes in the BCD. Fourth, we propose an alternative approach to content moderation that reduces toxicity while preserving the content of Tweets.
@@ 3.1 Removal of Toxic Content and Distortions of Online Discourse
We simulate the effects of more stringent content moderation by removing toxic Tweets based on varying toxicity thresholds from our data. We then analyze if removing toxic Tweets leads to distortions of online content as measured by the BCD. We will compare these content-moderation-induced changes in the semantic space to changes in the BCD if we instead remove the same number of Tweets from the data at random. Note that this analysis intentionally abstracts from any form of behavioral reaction on the side of the users, as we are interested in the direct effects of content moderation. This analysis should provide a realistic content moderation benchmark as the Perspectives API is used in real-world applications, and several studies of content moderation have used toxicity thresholds to delineate toxic content (e.g., Gehman et al., 2020; Han and Tsvetkov, 2020; Rieder and Skop, 2021; Hede et al., 2021; Jiménez Durán, 2022; Beknazar-Yuzbashev et al., 2022).
The results from this analysis are presented in Figure 1. The x-axis indicates the toxicity threshold above which we remove Tweets. We also report the share of Tweets that remain in the sample in parentheses below. The y-axis shows our measure of the distortion of the semantic space after the removal of toxic content relative to the original semantic space as measured by the BCD. It is immediately apparent that the removal of toxic content distortions to the semantic space (orange line). The findings indicate that distortions become more severe the lower we set the toxicity threshold. Importantly, removing Tweets from the data at random does not impact the BCD (blue line). This highlights that the increase in the BCD is not a mechanical consequence of a smaller sample of Tweets but rather driven by the changing composition of online discourse due to content moderation. We interpret this result as evidence that removing toxic distorts online content as measured by the embedding space.
@@
To benchmark the resulting distortions, we compare them to an approximation of the maximum possible distortion one could create by removing a specific number of Tweets. For this exercise, we remove the Tweets with the largest Euclidean distance from the centroid of the semantic space. Appendix Figure 1 shows the results from this exercise. We find that the removal of toxic content reaches around 20% of the maximum possible distortion. In Section 3.2, we provide additional benchmarks of the distortions relative to the direct removal of topics.
Robustness
We conduct several robustness exercises for this finding. First, we reproduce our finding using alternative embeddings based on the widely-used RoBerta and DeBerta models (see Figure C.2). Second, we base content moderation on the alternative Toxicity dimensions from the Perspectives API, the toxicity scores from the Detoxify classifiers (Hanu and Unitary team, 2020), and the Moderation API from OpenAI (see Figure C.3). Lastly, we repeat our analysis using samples of 1 million: 1) general Tweets (without filtering for political content), 2) German Tweets, and 3) Italian Tweets (see Figure C.4).
We discuss all of these robustness checks in greater detail in Appendix C. To summarize, none of these changes make any qualitative difference to our findings. Independent of the embedding model, the toxicity scores, or the sample, removing toxic content reduces the plurality of online discourse. Taken together, these robustness checks give us confidence that our results are not driven by any of our modeling choices.
3.2 What Drives Distortions to Online Content
We next aim to shed light on the factors that drive the content-moderation-induced distortions of the semantic space we have documented. First, we investigate the individual components of the BCD and show that both the mean as well as the variance of the embedding space get distorted by content moderation. Second, we show that the increase in the BCD is not driven by the use of toxic language itself. Third, we provide further evidence for the content-moderation-induced distortions of the topics of online debate.
@@ 3.2 What Drives Distortions to Online Content
As previously discussed, the BCD consists of two additive components (see eq. 1), the first of which captures distortions to the mean, while the latter captures distortions to the variance. We analyze to what extent each of these components contributes to the previously documented increases in the BCD. The results from this analysis are shown in Panel (a) of Figure 2. We observe that while both the first and the second component contribute to the increase in the BCD, the shift in the variance explains around 96% of the increase in the BCD, while the shift in the mean only explains 4%.
@@
Another interesting finding we can establish is that content moderation reduces the variance of the embedding space as measured by GVI ($\det\Sigma$) (see Panel (b) Figure 2). This subfigure plots the natural logarithm of the ratio of the GVI after content moderation relative to the original GVI. It is immediately apparent that the GVI decreases significantly with the removal of toxic content (orange line). Similar to our previous results, no change in the GVI occurs if content is removed at random (blue line). As the GVI captures the dispersion of content, the GVI is small if all Tweets are very similar and, therefore, close to each other in the semantic space. In contrast, if Tweets are widely dispersed in the semantic space, the variance will be large. The results suggest that content moderation appears to remove outliers in the semantic space. Removing such “outlier” content is arguably costly for the plurality of online speech.
Accounting for the Toxicity Dimension of the Embedding Space
The content moderation-induced changes in the semantic space could be driven by two factors. On the one hand, the toxic language itself could occupy a particular part of the semantic space. Removing toxic content would then mechanically increase the BCD, but this would not necessarily indicate a “real” distortions to online content. On the other hand, toxic Tweets might discuss specific issues and topics that are underrepresented online, independently of their toxicity.
We investigate this question by removing the toxicity component from the embedding space. To do so, we build on the literature on the debiasing of embeddings (e.g., Bolukbasi et al., 2016; Liang et al., 2020) and create orthogonal projections of the embeddings matrix $\mathbf{X}$ with respect to the Toxicity scores. Our baseline approach uses linear regressions of the following form:
$x_{d}=\alpha+\mathbf{Tox^{\prime}\beta}+\epsilon_{d}$
where $x_{d}\in\mathbf{X}$ is one of the $D$ embedding dimensions. $\mathbf{Tox}$ is a matrix containing 1000 indicators for the permilles of the toxicity distribution. [^14] We chose a non-parametric transformation of the toxicity scores to flexibly account for any potential non-linearities. As we show in a robustness check, the findings are almost identical if we instead residualize linearly with regard to the toxicity score. We estimate these regressions for all embedding dimensions $d\in D$ and replace the embedding dimensions with the regression residuals. This leaves us with a residualized matrix $\mathbf{\tilde{X}}$, which is orthogonal to the toxicity scores. In other words, we remove any variation in the embeddings of Tweets that can be explained by their toxicity. Any remaining variances in the embeddings should, therefore, capture differences in Tweets’ content independent of their toxicity.
We then repeat our previous analysis based on the residualized embedding matrix $\mathbf{\tilde{X}}$. Figure 3 visualizes the two different approaches to account for a Tweet’s toxicity. More specifically, we use regressions and either residualize the embeddings using permilles or the linear toxicity score.
We find that removing the toxicity component from the embedding space hardly makes any difference to our findings. The overall patterns are very close to our original results. The BCD sharply increases once toxic Tweets are removed from the data. This suggests that the observed changes in the BCD do not appear to be driven by the toxicity of the Tweets but rather by their content.
Figure 3: Controlling for Toxicity
Content-Moderation-Induced Topic Shifts
As the last piece of evidence on the content-moderation-induced shifts in online content, we analyze how the topic composition of Tweets is shifting with increased content moderation. Additionally, we provide evidence that, in line with our original motivation, the removal of topics is associated with increases in the BCD. The latter result also allows us to benchmark the content-moderation-induced increases in the BCD relative to the direct removal of topics from the embedding space.
In Panel (a) of Figure 4, we show the topics that would be affected most by content moderation. Perhaps unsurprisingly, we find that Tweets containing insults against political figures are most affected by content moderation. [^15] Note that even highly uncivil insults against public figures represent a form of opinion expression. Moreover, we also find that Tweets discussing the topics of “Racism”, “Trump”, or the “Black Lives Matter (BLM)” movement are heavily affected by content moderation, thereby distorting the topic composition of online debate. [^16] The finding on the removal of discussions of racism aligns well with the work of Lee et al. (2024). It is also important to highlight that our results do not hinge on shifts in any one topic, i.e., the results are very virtually identical if we do not consider Tweets containing insults (see Appendix Figure C.5). Further, we can show that content moderation distorts the semantic space even within topics (unreported).
Figure 4: Content Moderation and Topic Shifts
(a) Topics Shifts
(b) BCD and Removal of Topics
Notes: The figure visualizes the effect of content moderation on changes in topics as generated by the Top2Vec algorithm. Panel (a) shows which topics are most heavily affected by content moderation. Panel (b) shows that removing 1, 2, 3, 4, or 5 topics from the data at random leads to increases in the BCD. The percentages in parentheses on the x-axis represent the proportion of Tweets retained relative to the original sample size.
In Panel (b), we lastly show that removing topics leads to increasing changes in the BCD. For this figure, we calculate the BCD for the cases in which we remove 1, 2, 3, 4, or 5 topics from the data at random. For this exercise, we implement 25 random draws of the indicated number of topics. We then remove Tweets from the selected topics from the data and calculate the BCD. The figure then reports the average BCD resulting from the 25 draws. We also report the average share of Tweets that remain in the data after the removal of topics in brackets on the x-axis.
We find that the BCD increases progressively with the removal of additional topics. This exercise also allows us to benchmark the content-moderation-induced distortions we documented in Figure 1. We find that removing all Tweets with a toxicity score above 0.8 from the data is akin to removing four topics from the data at random. Both of these interventions result in a BCD of ca. 0.7. It is also worth highlighting that removing four deletes twice as many Tweets from the data (6%) as content moderation with a threshold of 0.8 (3%). This suggests that the per-Tweet distortions of the semantic space are more severe when removing content based on toxicity.
Taken stock, the previous exercises suggest that the reduction in BCD are not driven by the toxic language itself, but rather the fact that removing Tweets with toxic language as a byproduct 1) shifts the mean of the embeddings space, 2) reduces the variance of the semantic space, and 3) removes topics from the online debate. In the last part of the paper, we propose an alternative approach to mitigate some of the content-moderation-induced distortions.
3.3 An Alternative Approach to Content Moderation

### v2

@@ 2.2 Measuring Distortions to Online Content
As described in the introduction, our measure of content distortions is based on the idea of the semantic space. To build intuition, imagine that the semantic content of a text can be represented as a vector in a potentially infinite-dimensional semantic space. In this space, texts that talk about the same issue or hold the same opinion are close together, while texts about other issues or diverging opinions are far apart. Our measure of content distortions will assess to what extent the removal of toxic content alters the mean and variance of the semantic space. The measure will be small if the semantic space after the removal of toxic content is very similar to the original semantic space prior to content moderation. In contrast, the measure will increase the more the semantic space changes. Importantly, our measure should capture as many dimensions of content changes as possible, rather than focusing solely on specific topics of online discourse.
Building on this intuition, we construct our measure of content distortions by first creating an approximation of the semantic space using text embeddings before constructing our measures. We describe each of these steps in the following. First, we create embeddings for each of the Tweets in our data using the BERTweet model (Nguyen et al., 2020). [^9] We also provide robustness checks for embeddings created by the RoBerta, DeBerta, and DistilBert-Base-Multilingual-Cased-V2 models. The BERTweet model, trained on a large English-language Twitter corpus, transforms the text of each Tweet into a 768-dimensional vector. In this way, the BERTweet model generates an approximation of the semantic space. [^10] Additional details on the BERTweet model can be found in Section B.1.
@@
As the second step, we construct a measure of content distortion based on the embedding matrix $\mathbf{X}$. Our measure is based on the Bhattacharyya distance (BCD), a commonly used metric for comparing probability distributions. As the embeddings are approximately normally distributed (see Appendix Figure B.1), we use the formula for Bhattacharyya distance for multivariate normal distributions (e.g., Abou-Moustafa and Ferrie, 2012). For the case of two multivariate normal distributions $\mathcal{N}_{1}(\mu_{1},\Sigma_{1})$ and $\mathcal{N}_{2}(\mu_{2},\Sigma_{2})$, the BCD is defined as:
@@
where $\Sigma=\frac{\Sigma_{1}+\Sigma_{2}}{2}$. The Bhattacharyya consists of two additive terms. The first term is the squared Mahalanobis distance (Mahalanobis, 1936). It measures the inverse-variance-weighted distortions to the mean of the multivariate normal distribution, while the second term measures distortions to the variance. The $\det\Sigma$ is the so-called generalized variance index (GVI) (Wilks, 1932), which provides a multivariate extension of the standard statistical variance measure based on the mean squared deviation.
To calculate the BCD, we calculate both the means ($\mu$) and the variance-covariance matrix ($\Sigma$) of the embedding matrix $\mathbf{X}$. [^11] As the calculation of the dot product $XX^{\prime}$ is computationally unfeasible, we instead use a maximum likelihood estimator of the empirical covariance matrix using the procedure implemented in scikit learn. In tests on a subsample of the data, we confirmed that the approximation is very close to the analytical solution of the variance-covariance matrix. The diagonal elements of the variance-covariance matrix capture the dispersion of Tweets along the embedding dimensions. Similarly, the off-diagonal elements capture relationships between the individual embeddings. [^12] The individual entries in $\Sigma(\mathbf{X})$ are relatively small, which can lead to integer underflow when calculating the GVI. To avoid this issue, we multiply the embedding matrix $\mathbf{X}$ by 100 before calculating the BCD. Given that this equally affects all components of the BCD, this has no bearing on our results.
In this way, BCD provides a unidimensional measure that summarizes the overall distortions of the embedding matrix $\mathbf{X}$. The BCD has several properties that make it a desirable measure for our application:
@@
    Another key advantage of the BCD is its computational tractability in high-dimensional settings. The BCD allows us to approximate the distortions of the embedding space in a computationally feasible manner, a crucial aspect given the expansive nature of online platforms and our data. [^13] Many other alternative methods, such as the ones that involve the computation of convex hulls, become computationally infeasible in high-dimensional settings.
Advantages over Potential Alternative Metrics
Readers familiar with the broader natural language processing literature may wonder why we propose a new measure to quantify content distortions, rather than relying on more commonly used approaches. The two most plausible candidates could be to calculate the cosine similarity of the embedding vectors or use topic models to analyse how the topic distribution has shifted.
While these techniques are very useful and we will use a topic model later in our analyses, our approach has several advantages compared to these alternative measures. First, our measure is computationally significantly cheaper to construct from the embedding space than cosine-based measures, which is particularly relevant given the large size of social media platforms. In comparison, the required number of calculations and memory grow quadratically in the number of observations for the cosine similarity. Secondly, both cosine similarity and topic models may fail to detect semantic changes in online content, despite the social media content having undergone significant shifts. In Appendix C, we show that the average similarity of tweets remains virtually unchanged when toxic content is removed. Similarly, topic models do not account for the differences in opinions expressed within a topic (e.g., pro- and anti-immigration tweets). It would therefore be possible to remove one side from the political debate without necessarily shifting the topic composition. For example, if one political side accounted for half of each online topic, we could remove this side completely from each topic without altering the underlying distribution of topics. In contrast, our approach provides a holistic measure for the total extent of semantic shifts.
@@ 3 Results
In the following, we use BCD to establish three sets of results. First, we analyze the extent to which online content is shifted when we remove highly toxic Tweets from the data. This analysis allows us to simulate the effect of more stringent content moderation. Second, we benchmark the magnitude of these shifts and show that the content-moderation-induced shifts are comparable to the removal of entire topics from the online debate. Third, we demonstrate that content moderation-induced distortions can be reduced by rephrasing highly toxic Tweets in a way that removes the toxic language while maintaining the original content. The last finding suggests that the documented shifts in online content are not exclusively driven by toxic language, but rather stem from the removal of specific topics from the online debate.
@@ 3.1 Removal of Toxic Content and Distortions of Online Discourse
In our initial analysis, we simulate the effects of more stringent content moderation by removing toxic Tweets based on varying toxicity thresholds from our data. This analysis provides a realistic content moderation benchmark as the Perspectives API is used in real-world applications, and several studies of content moderation have used toxicity thresholds to delineate toxic content (e.g., Gehman et al., 2020; Han and Tsvetkov, 2020; Rieder and Skop, 2021; Hede et al., 2021; Jiménez Durán, 2022; Beknazar-Yuzbashev et al., 2022). We then use BCD to analyze whether removing toxic Tweets leads to distortions of online content. We compare these content-moderation-induced changes to a baseline by removing the same number of Tweets from the data at random.
Note that this analysis intentionally abstracts from any behavioral user reaction, as we are interested in the direct effects of content moderation. For example, the removal of toxic content could allow non-toxic users to more freely express their opinions, or toxic users could be deterred from using toxic language. In our analysis, we aim to isolate the first-round effects of content moderation; however, our measure would also be well-suited to characterize any behavioral reactions.
The results from our first analysis are presented in Figure 1. The x-axis indicates the toxicity threshold above which we remove Tweets. We also report the share of Tweets that remain in the sample in parentheses below. The y-axis shows our measure of the distortion of the semantic space after the removal of toxic content relative to the original semantic space as measured by the BCD. It is immediately apparent that the removal of toxic content leads to distortions to the semantic space (orange line). Intuitively, the findings indicate that changes become more severe the lower we set the toxicity threshold. Importantly, removing Tweets from the data at random does not impact the BCD (blue line). This highlights that the increase in the BCD is not a mechanical consequence of a smaller sample of Tweets but rather is driven by the changing composition of online discourse due to content moderation. This finding suggests that the removal of toxic online content leads to significant shifts in the embedding space.
@@
Benchmarking the Magnitude of Distortions
Next, we provide two benchmarks to better understand the magnitude of the content-moderation-induced alterations of the semantic space: 1) an upper bound for BCD, and 2) the effect of topic-based removal.
First, we compare changes in the BCD when removing Tweets based on their toxicity scores versus removing an equal number of Tweets with the greatest Euclidean distance from the centroid of all embedded Tweets. Given the definition of Bhattacharyya distance, removing Tweets furthest from the distribution’s center approximates the upper bound of BCD changes possible by removing a specific number of Tweets. Figure 2 shows the changes in BCD when removing Tweets with toxicity scores above a certain threshold and compares this to removing an equal number of Tweets based on their distance from the centroid. The results indicate that removing toxic Tweets increases BCD by approximately 20% of the maximum possible increase from removing Tweets with a large Euclidean distance.
Figure 2: Benchmarking BCD
Second, we provide an additional benchmark of the distortions relative to the direct removal of topics. In Panel (a) of Figure 3, we analyze how much the semantic space changes if we directly delete Tweets discussing specific topics from the data. For this figure, we calculate the BCD for cases in which we randomly remove 1, 2, 3, 4, or 5 topics from the data. In each case, we implement 25 random draws of the indicated number of topics. We then remove Tweets from the selected topics from the data and calculate the BCD. The figure then reports the average BCD resulting from the 25 draws. We also report the average share of Tweets that remain in the data after the removal of topics in brackets on the x-axis.
Intuitively, we observe that the BCD increases progressively as additional topics are removed. This exercise also allows us to provide another benchmark for the content-moderation-induced distortions we documented in Figure 1. We find that removing all Tweets with a toxicity score above 0.8 from the data is akin to removing four topics from the data at random. Both of these interventions result in a BCD of approximately 0.7. However, it is also worth highlighting that removing four topics deletes twice as many Tweets from the data (6%) as content moderation with a threshold of 0.8 (3%). This suggests that the per-Tweet distortions of the semantic space are more severe when content is removed based on toxicity.
In Panel (b) of Figure 3, we additionally show the topics that are affected most by content moderation. Perhaps unsurprisingly, we find that Tweets containing insults against political figures are most affected by content moderation. [^14] Note that even highly uncivil insults against public figures represent a form of opinion expression. Moreover, we also find that Tweets discussing the topics of “Racism”, “Trump”, or the “Black Lives Matter (BLM)” movement are heavily affected by content moderation, thereby distorting the topic composition of online debate. [^15] The finding on the removal of discussions of racism aligns well with the work of Lee et al. (2024). It is also important to highlight that our results do not hinge on shifts in any one topic, i.e., the results are virtually identical if we do not consider Tweets containing insults (see Appendix Figure C.6). Furthermore, we can demonstrate that content moderation distorts the semantic space, even within specific topics (unreported).
Figure 3: Content Moderation and Topic Shifts
(a) BCD and Removal of Topics
(b) Topics Shifts
Notes: The figure visualizes the effect of content moderation on changes in topics as generated by the Top2Vec algorithm. Panel (a) shows that removing 1, 2, 3, 4, or 5 topics from the data at random leads to increases in the BCD. Panel (b) shows which topics are most heavily affected by content moderation. The percentages in parentheses on the x-axis represent the proportion of Tweets retained relative to the original sample size.
@@
As the next step, we develop a better understanding of the statistical moments (mean and variance) that underlie increases in the BCD. As previously discussed, the BCD consists of two additive components (see eq. 1), the first of which captures distortions to the mean, while the latter captures distortions to the variance. We analyze the extent to which each of these components contributes to the previously documented increases in the BCD. The results from this analysis are shown in Panel (a) of Figure 4. We observe that while both the first and second components contribute to the increase in the BCD, the shift in variance explains approximately 96% of the increase in the BCD, whereas the shift in the mean only explains 4%.
@@
Another interesting finding we can establish is that content moderation reduces the variance of the embedding space as measured by GVI ($\det\Sigma$) (see Panel (b) Figure 4). This subfigure plots the natural logarithm of the ratio of the GVI after content moderation relative to the original GVI. It is immediately apparent that the GVI decreases significantly with the removal of toxic content (orange line). Similar to our previous results, no change in the GVI occurs if content is removed at random (blue line). As the GVI captures the dispersion of content, the GVI is small if all Tweets are very similar and, therefore, close to each other in the semantic space. In contrast, if Tweets are widely dispersed in the semantic space, the variance will be large. The results suggest that content moderation appears to remove outliers in the semantic space. Removing such “outlier” content is arguably costly for the plurality of online speech as they represent positions that are expressed less frequently online.
Robustness
We conduct several robustness checks to verify our findings. First, we reproduce our finding using alternative embeddings based on the widely used RoBERTa, DeBERTa, and DistilBert-Base-Multilingual-Cased-V2 models (see Figure C.2). Second, we base content moderation on the alternative Toxicity dimensions from the Perspectives API, the toxicity scores from the Detoxify classifiers (Hanu and Unitary team, 2020), and the Moderation API from OpenAI (see Figure C.3). Third, we show that our results are very similar when we weight Tweets by their user engagement, as proxied by the number of Retweets. Lastly, we repeat our analysis using samples of 1 million: 1) general Tweets (without filtering for political content), 2) German Tweets, and 3) Italian Tweets (see Figure C.5).
We discuss all of these robustness checks in greater detail in Appendix C. To summarize, none of these changes makes any qualitative difference to our findings. Independent of the embedding model, the toxicity scores, or the sample, removing toxic content reduces the plurality of online discourse. Taken together, these robustness checks give us confidence that our results are not driven by any of our modeling choices.
3.2 What Drives Distortions to Online Content
The previously documented distortions may arise for two reasons. On the one hand, they could reflect a mechanical shift in the semantic space caused by the removal of toxic language. Abstracting from debates over what constitutes toxicity, such shifts would arguably not be costly, since they merely reflect the removal of content already deemed unacceptable for the platform. On the other hand, toxic Tweets might discuss specific issues and topics that are underrepresented online using inflammatory language. In this case, removing such Tweets would distort the online debate beyond the elimination of the toxic language itself.
We evaluate these competing hypotheses using two complementary approaches. First, we use large language models (LLMs) to rephrase Tweets, stripping away toxic language while preserving the underlying message. We show that replacing Tweets with their rephrased counterparts (instead of removing them entirely) mitigates some of the distortions to the semantic space. This result highlights that the distortions of the semantic space are not entirely driven by the removal of toxic language. Furthermore, this exercise demonstrates the potential of our measure to benchmark different content moderation strategies against one another.
Second, building on the literature on debiasing text embeddings (e.g., Bolukbasi et al., 2016), we demonstrate that moderation-induced distortions are not attributable to a specific spatial positioning of toxic language in the embedding space, but rather to changes in online content. Specifically, we construct projections of the embedding space orthogonal to toxicity scores (i.e., embeddings that assign the same vector representation to Tweets regardless of their toxicity). We find that content moderation still induces distortions in this orthogonalized embedding space, suggesting that the distortions are not merely an artifact of removing toxic language. Furthermore, we show that removing toxic Tweets in the orthogonalized embedding space, as opposed to removing their rephrased counterparts in the original embedding space, produces nearly identical BCDs.
Rephrasing Tweets